# Poincaré inequality on the monotone sets revisited

### v1

@@ 2 Preliminaries and Proofs
$\begin{split}a&:=\mu^{n}(A)=\frac{1}{2^{n}}\sum_{x\in V^{n}}\mathbbm{1}_{A}(x)=\frac{1}{2}\cdot\frac{1}{2^{n-1}}\left(\sum_{(x^{\prime},1)\in V_{1}}\mathbbm{1}_{A}(x^{\prime},1)+\sum_{(x^{\prime},0)\in V_{0}}\mathbbm{1}_{A}(x^{\prime},0)\right)\\ &=\frac{1}{2}\cdot\frac{1}{2^{n-1}}\left(\sum_{x^{\prime}\in V^{n-1}}\mathbbm{1}_{A_{1}}(x^{\prime})+\sum_{x^{\prime}\in V^{n-1}}\mathbbm{1}_{A_{0}}(x^{\prime})\right)=\frac{1}{2}\mu^{n-1}(A_{1})+\frac{1}{2}\mu^{n-1}(A_{0})=\frac{a_{0}+a_{1}}{2}.\end{split}$ (5)
@@
Fact 2.1.
If $A$ is an increasing set, then $A_{0}\subset A_{1}$.
Proof.
For any $x^{\prime}\in A_{0}$, we shall see that $x^{\prime}\in A_{1}$. It is easy to see because $(x^{\prime},0)\leq(x^{\prime},1)$ and $A$ is increasing. ∎
By applying conditional expectation expansion to $n$th variable of restricted Dirichlet forms, we can readily reduce the Dirichlet form of the original function to that of the restricted functions.
@@
We use $x^{\prime}$ to denote the first $(n-1)$-coordinates of $x$. Note that
$\begin{split}\mathcal{E}_{A}(f)&=\frac{1}{4}\cdot\sum\limits_{i=1}^{n-1}\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus i}))^{2}\cdot\mathbbm{1}_{\{x^{\oplus i}\in A\}}\right]+\frac{1}{4}\cdot\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus n}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1-x_{n})\in A\}}\right]\\ &=\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},1)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},0)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},0)\in A\}}\right]+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1)\in A\}}\right]\\ &=\frac{a_{1}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{1}\}}\right]+\frac{a_{0}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{0}\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right]+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right],\end{split}$
@@
Similarly, we can reduce the variance of the original function to that of the restricted functions.
@@
$\begin{split}\mu&=\frac{1}{|A|}\sum_{x\in A}f(x)=\frac{1}{|A|}\sum_{x^{\prime}\in A_{1}}f(x^{\prime},1)+\frac{1}{|A|}\sum_{x^{\prime}\in A_{0}}f(x^{\prime},0)=\frac{|A_{1}|}{|A|}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]+\frac{|A_{0}|}{|A|}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\\ &=\frac{a_{1}}{2a}\mu_{1}+\frac{a_{0}}{2a}\mu_{0}.\end{split}$ (8)
Thus
$\begin{split}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]&=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[\left|f_{1}-\frac{a_{1}}{2a}\mu_{1}-\frac{a_{0}}{2a}\mu_{0}\right|^{2}\right]=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[\left|f_{1}-\mu_{1}+\mu_{1}-\frac{a_{1}}{2a}\mu_{1}-\frac{a_{0}}{2a}\mu_{0}\right|^{2}\right]\\ &=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[\left|f_{1}-\mu_{1}+\frac{a_{0}}{2a}(\mu_{1}-\mu_{0})\right|^{2}\right]\\ &=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[\left|f_{1}-\mu_{1}\right|^{2}\right]+2\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}-\mu_{1})\cdot\frac{a_{0}}{2a}(\mu_{1}-\mu_{0})\right]+\frac{a_{0}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2}\\ &={\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2}.\end{split}$
Similarly, we have
@@
$\begin{split}{\rm Var}_{A}[f]&=\frac{a_{1}}{2a}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]+\frac{a_{0}}{2a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{1}a_{0}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}a_{1}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}^{2}+a_{0}a_{1}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot(\mu_{1}-\mu_{0})^{2}\end{split}$
3 An inductive proof of Theorem 1.3
Our goal is to prove that for all $f:A\to\mathbb{R}$,
@@
where $c>0$ is a universal constant (we can take $c=0.5$). It is easy to check that the theorem holds for $n=1$. Let $n\geq 2$, and suppose the statement of the theorem holds when $n$ is replaced by $n-1$. Using Lemma 2.2 to split the Dirichlet form of the original function to two restricted functions:
@@
$\begin{split}\mathcal{E}_{A}(f)&=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\geq\frac{a_{1}}{2a}\cdot(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}(c-c\sqrt{1-a}){\rm Var}_{A}[f].\end{split}$ (12)
@@
${\rm Var}_{A}[f]=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}^{2}+a_{0}a_{1}^{2}}{8a^{3}}\left(\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\right)^{2}.$ (13)
@@
$\begin{split}&\frac{a_{1}}{2a}\cdot(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}\frac{a_{1}}{2a}\cdot(c-c\sqrt{1-a}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot(c-c\sqrt{1-a}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot(c-c\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (14)
@@
$\begin{split}&ca_{1}\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+ca_{0}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{2}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}c\frac{a_{1}a_{0}}{a_{0}+a_{1}}(1-\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (15)
@@
$\begin{split}&\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{0}(x^{\prime})-\mu_{0}-(f_{1}(x^{\prime})-\mu_{1})+\mu_{0}-\mu_{1}|^{2}\right]\\ &=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{0}-\mu_{0}|^{2}\right]+\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{1}-\mu_{1}|^{2}\right]+2\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]+(\mu_{0}-\mu_{1})^{2}+2(\mu_{0}-\mu_{1})(\mu_{1}-\alpha).\end{split}$ (17)
@@
$\begin{split}&\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\left(C_{0}+\frac{a_{0}}{2}\right)\cdot{\rm Var}_{A_{0}}[f_{0}]+a_{0}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &=\frac{1}{|A_{0}|}\sum\limits_{x^{\prime}\in A_{0}}\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)|\mu_{1}-f_{1}(x^{\prime})|^{2}+\left(C_{0}+\frac{a_{0}}{2}\right)|f_{0}(x^{\prime})-\mu_{0}|^{2}+a_{0}(f_{0}(x^{\prime})-\mu_{0})(\mu_{1}-f_{1}(x^{\prime}))\\ &=\frac{1}{|A_{0}|}\sum_{x^{\prime}\in A_{0}}\vec{b}_{x^{\prime}}^{\top}G_{x^{\prime}}\vec{b}_{x^{\prime}},\end{split}$ (19)
@@
$\vec{b}_{x^{\prime}}=\begin{bmatrix}f_{0}(x^{\prime})-\mu_{0}\\ \mu_{1}-f_{1}(x^{\prime})\end{bmatrix},\\ G_{x^{\prime}}=\begin{bmatrix}a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{0}}+1}{2}&\frac{a_{0}}{2}\\ \frac{a_{0}}{2}&a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{1}}+1}{2}.\end{bmatrix}\$
@@
Let $s:=\sqrt{1-a_{0}}$ and $t:=\sqrt{1-a_{1}}$, then $0\leq t\leq s\leq 1$, $21-a=(s^{2}+t^{2})/$ 2 and

### v2

@@ 3.1 Preliminaries
$\begin{split}a:=\mu^{n}(A)=\frac{1}{2}\mu^{n-1}(A_{1})+\frac{1}{2}\mu^{n-1}(A_{0})=\frac{a_{0}+a_{1}}{2}.\end{split}$ (23)
@@
By expanding the Dirichlet form with respect to the conditional expectation on the $n$th variable, we can readily express the Dirichlet form of the original function in terms of those of its restricted versions.
@@
We use $x^{\prime}$ to denote the first (n1)(n-1)-coordinates of $x$. Note that
$\begin{split}\mathcal{E}_{A}(f)&=\frac{1}{4}\cdot\sum\limits_{i=1}^{n-1}\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus i}))^{2}\cdot\mathbbm{1}_{\{x^{\oplus i}\in A\}}\right]+\frac{1}{4}\cdot\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus n}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1-x_{n})\in A\}}\right]\\ &=\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},1)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},0)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},0)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1)\in A\}}\right]\\ &=\frac{a_{1}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{1}\}}\right]\\ &+\frac{a_{0}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{0}\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right]+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right],\end{split}$
@@
Similarly, by conditioning on the $n$th coordinate, the variance of the original function can be expressed in terms of the variances of its restricted functions.
@@
$\mu=\frac{1}{|A|}\sum_{x\in A}f(x)=\frac{|A_{1}|}{|A|}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]+\frac{|A_{0}|}{|A|}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]=\frac{a_{1}}{2a}\mu_{1}+\frac{a_{0}}{2a}\mu_{0}.$ (26)
Thus by the formula $\mathbb{E}[(f-b)^{2}]={\rm Var}[f]+(\mathbb{E}[f]-b)^{2}$, we have that
$\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]={\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2},$
and
@@
$\begin{split}{\rm Var}_{A}[f]&=\frac{a_{1}}{2a}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]+\frac{a_{0}}{2a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{1}a_{0}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}a_{1}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot(\mu_{1}-\mu_{0})^{2}\end{split}$
3.2 Induction part
Our goal is to prove that for every function $f:A\to\mathbb{R}$,
@@ 3.2 Induction part
where $c>0$ is a universal constant (we may take $c=0.5$). It is easy to verify that the theorem holds for $n=1$. Let $n\geq 2$, and suppose the statement of the theorem holds when $n$ is replaced by n1n-1. Applying Lemma 3.1 to split the Dirichlet form of the original function into two restricted functions:
@@
$\begin{split}\mathcal{E}_{A}(f)&=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\geq\frac{a_{1}}{2a}(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{4a}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}(c-c\sqrt{1-a}){\rm Var}_{A}[f].\end{split}$ (30)
@@
${\rm Var}_{A}[f]=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\left(\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\right)^{2}.$ (31)
@@
$\begin{split}&\frac{a_{1}}{2a}\cdot(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}\frac{a_{1}}{2a}(c-c\sqrt{1-a}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}(c-c\sqrt{1-a}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}(c-c\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (32)
@@
$\begin{split}&ca_{1}\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+ca_{0}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]\\ &+\frac{a_{0}}{2}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\underset{?}{\geq}c\frac{a_{1}a_{0}}{a_{0}+a_{1}}(1-\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (33)
@@
$\begin{split}&\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{0}(x^{\prime})-\mu_{0}-(f_{1}(x^{\prime})-\mu_{1})+\mu_{0}-\mu_{1}|^{2}\right]\\ &=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{0}-\mu_{0}|^{2}\right]+\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{1}-\mu_{1}|^{2}\right]+2\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &+(\mu_{0}-\mu_{1})^{2}+2(\mu_{0}-\mu_{1})(\mu_{1}-\alpha).\end{split}$ (35)
@@
$\begin{split}&\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\left(C_{0}+\frac{a_{0}}{2}\right)\cdot{\rm Var}_{A_{0}}[f_{0}]+a_{0}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &=\frac{1}{|A_{0}|}\sum\limits_{x^{\prime}\in A_{0}}\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)|\mu_{1}-f_{1}(x^{\prime})|^{2}+\left(C_{0}+\frac{a_{0}}{2}\right)|f_{0}(x^{\prime})-\mu_{0}|^{2}\\ &+a_{0}(f_{0}(x^{\prime})-\mu_{0})(\mu_{1}-f_{1}(x^{\prime}))\\ &=\frac{1}{|A_{0}|}\sum_{x^{\prime}\in A_{0}}\vec{b}_{x^{\prime}}^{\top}G_{x^{\prime}}\vec{b}_{x^{\prime}},\end{split}$ (37)
@@
$\vec{b}_{x^{\prime}}=\begin{bmatrix}f_{0}(x^{\prime})-\mu_{0}\\ \mu_{1}-f_{1}(x^{\prime})\end{bmatrix},\\ G_{x^{\prime}}=\begin{bmatrix}a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{0}}+1}{2}&\frac{a_{0}}{2}\\ \frac{a_{0}}{2}&a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{1}}+1}{2}\end{bmatrix}$.\\
@@
Let $s:=\sqrt{1-a_{0}}$ and $t:=\sqrt{1-a_{1}}$, then $0\leq t\leq s\leq 1$, $1-a=(s^{2}+t^{2})/2$ and